# A Safe Preference Learning Approach for Personalization with Applications to Autonomous Vehicles

### v3

@@ III-B Weighted Signal Temporal Logic (WSTL)
[…]
@@
In the WSTL definition of [12], weights are pre-determined positive real values. In this work, we use an extension to WSTL that we call Parametric Weighted Signal Temporal Logic (PWSTL) in which some of the weights are unknown parameters and the remaining weights are given constants (cf., [31]). We denote the set of unknown parameters as $\mathcal{W}$ and denote PWSTL formulas as $\phi_{\mathcal{W}}$, where we omit the known weights with slight abuse of notation since for most of the results in the paper $\mathcal{W}$ is the entire weight set. A PWSTL formula results in a WSTL formula $\phi_{\mathcal{W}=w}$ with the valuation $w$ of the parameters.
@@ IV-A An Optimization Reformulation
where $\mathbb{1}(w)$ is the indicator function which takes $\mathbb{1}(w)=1$ when the subscripted condition is satisfied and takes $\mathbb{1}(w)=0$ otherwise. [^3] Note that $\mathcal{D}$ is an open set, but the objective function takes only finitely many values, hence it always has a minimum. Therefore, searching for $\arg\min$ is valid.
@@
We note that for certain STL formulas $\phi_{\rm STL}$, the corresponding WSTL robustness metric satisfies $r(s,\phi_{\mathcal{W}=w})\leq 0$ for all signals $s$ and all weights $w\in\mathbb{R}_{+}^{n}$. In this case, the problem of learning weights only from rule-satisfying signals is not meaningful since these signals will have $r(s,\phi_{\mathcal{W}=w})=0$ for any weight $w$; and satisfaction cannot be deduced when $r(s,\phi_{\mathcal{W}=w})=0$. While such formulas can be uncommon in certain domains, we find them to be common in driving scenarios when the specification involves coming to a full stop or a Boolean indicator for pedestrians or traffic lights. Hence, we discuss a workaround to enable preference learning among satisfying signals for a class of such formulas but similar workarounds can be devised for other cases too. Consider a formula of the form $\phi=\phi_{1}\wedge\square(f(s)\geq 0\;\wedge\;-f(s)\geq 0)$, second part of which essentially represents an equality $f(s)=0$. To enable preference learning in this case, we append $s$ with a new coordinate $b$ such that $b(t)=\infty$ if $f(s(t))=0$ and $b(t)=-\infty$ otherwise; and replace the formula with $\phi^{\prime}=\phi_{1}\wedge\square b$. With this transformation, if for all $t\in[0,t_{final}]$, $b(t)=\infty$, the WSTL robustness of ϕsuperscriptitalic-$\phi^{\prime}$ is determined by $\phi_{1}$; otherwise the WSTL robustness of ϕsuperscriptitalic-$\phi^{\prime}$ becomes $\infty$- ∞, which indicates a violation of $\phi$. Applying our method to ϕsuperscriptitalic-$\phi^{\prime}$ allows us to rank rule-satisfying signals.
@@ V-A Baseline Methods
where $\psi(\cdot)$ again represents the feature vector described earlier. Then, we solve for the weights $v$ to maximize the log-likelihood as follows:
$v^{*}=\arg\min-\sum_{i=1}^{P}{\log(P_{v}(s_{i}^{+},s_{i}^{-}))}.$ (4)
@@ V-B Comparison of Solution Approaches
Some representative results are shown in Figures 2(a) and 2(b). The average performance of all methods for all users and splits is shown in Table I. Random sampling gives competitive results with the baseline methods in the stop sign scenario, and outperforms all methods in the pedestrian. Although the gradient-based method gives equally good results as random sampling for the stop sign scenario, it is much slower. Indeed, it was too slow to converge for the pedestrian scenario, due to formula complexity, and is skipped. We conclude that simple random sampling effectively identifies promising weight valuations that improve the traditional STL accuracy, and give comparable results to other methods.

### v4

@@ III-B Weighted Signal Temporal Logic (WSTL)
Let $\tilde{r}:\mathcal{S}\times\mathcal{F}\times\mathbb{T}\to\mathbb{R}$ be a quantitative semantics. For a WSTL formula $\phi$, let $\phi_{\rm STL}$ be the STL formula obtained by removing the weights in $\phi$. If $\text{sign}(\rho(s,\phi_{\rm STL},t))=\text{sign}(\tilde{r}(s,\phi,t))$ for all $(s,\phi,t)\in\mathcal{S}\times\mathcal{F}\times\mathbb{T}$ (i.e., $\tilde{r}$ is sign-consistent), then $\tilde{r}$ is sound.
@@
In the WSTL definition of [12], weights are pre-determined positive real values. In this work, we use an extension to WSTL that we call Parametric Weighted Signal Temporal Logic (PWSTL) in which some of the weights are unknown parameters and the remaining weights are given constants (cf., [31]). We denote the set of unknown parameters as $\mathcal{W}$ and denote PWSTL formulas as $\phi_{\mathcal{W}}$, where we omit the known weights with slight abuse of notation since for most of the results in the letter, $\mathcal{W}$ is the entire weight set. A PWSTL formula results in a WSTL formula $\phi_{\mathcal{W}=w}$ with the valuation $w$ of the parameters.
@@ IV-A An Optimization Reformulation
where $\mathbb{1}(w)$ is the indicator function which takes $\mathbb{1}(w)=1$ when the subscripted condition is satisfied and takes $\mathbb{1}(w)=0$ otherwise. [^3] Since the objective function takes only finitely many values, it always has a minimum. Therefore, searching for $\arg\min$ is valid.
@@
We note that for certain STL formulas $\phi_{\rm STL}$, the corresponding WSTL robustness metric satisfies $r(s,\phi_{\mathcal{W}=w})\leq 0$ for all signals $s$ and all weights $w\in\mathbb{R}_{+}^{n}$. In this case, the problem of learning weights only from rule-satisfying signals is not meaningful since these signals will have $r(s,\phi_{\mathcal{W}=w})=0$ for any weight $w$; and satisfaction cannot be deduced when $r(s,\phi_{\mathcal{W}=w})=0$. While such formulas can be uncommon in certain domains, we find them to be common in driving scenarios when the specification involves coming to a full stop or a Boolean indicator for pedestrians or traffic lights. Hence, we discuss a workaround to enable preference learning among satisfying signals for a class of such formulas but similar workarounds can be devised for other cases too. Consider a formula of the form $\phi=\phi_{1}\wedge\square(f(s)\geq 0\;\wedge\;-f(s)\geq 0)$, second part of which essentially represents an equality $f(s)=0$. To enable preference learning in this case, we append $s$ with a new coordinate $b$ such that $b(t)=\infty$ if $f(s(t))=0$ and $b(t)=-\infty$ otherwise; and replace the formula with $\phi^{\prime}=\phi_{1}\wedge\square b$. With this transformation, if for all $t\in[0,t_{final}]$, $b(t)=\infty$, the WSTL robustness of $\phi^{\prime}$ is determined by $\phi_{1}$; otherwise the WSTL robustness of $\phi^{\prime}$ becomes $\infty$- ∞, which indicates a violation of $\phi$. Applying our method to $\phi^{\prime}$ allows us to rank rule-satisfying signals.
@@ V-A Baseline Methods
where $\psi(\cdot)$ again represents the feature vector described earlier. Then, we solve for weights $v$ to maximize the log-likelihood as follows:
$v^{*}\in\arg\min-\sum_{i=1}^{P}{\log(P_{v}(s_{i}^{+},s_{i}^{-}))}.$ (4)
@@ V-B Comparison of Solution Approaches
Some representative results are shown in Figures 3(a) and 3(b). The average performance of all methods for all users and splits is shown in Table I. Random sampling gives competitive results with the baseline methods in the stop sign scenario, and outperforms all methods in the pedestrian. Although the gradient-based method gives equally good results as random sampling for the stop sign scenario, it is much slower. Indeed, it was too slow to converge for the pedestrian scenario, due to formula complexity, and is skipped. We conclude that simple random sampling effectively identifies promising weight valuations that improve the traditional STL accuracy, and give comparable results to other methods.